Suppose that the total cost equation (TC) for a monopolist is given by
TC = 500 + 20Q2
Let the demand equation be given by
P = 400 20Q
What are the profit-maximizing price and quantity?
1
Expert's answer
2020-11-23T10:21:34-0500

The profit-maximizing quantity can be found at MR = MC:

"MR = TR'(Q) = 400 - 40Q,"

"MC = TC'(Q) = 40Q,"

"400 - 40Q = 40Q,"

Q = 5 units,

P = 400 - 20×5 = 300.